OpenStudy (anonymous):

this one is weird... (-40^2)(30^6) over -50^6

OpenStudy (turingtest):

I would factor everything and see what cancels I guess

OpenStudy (anonymous):

i still dont know how to solve it :(

OpenStudy (mertsj):

\[\frac{4^2 \times 10^2\times3^6 \times 10^6}{5^6 \times 10^6}\]

OpenStudy (mertsj):

The 10^6 cancels so we have :\[\frac{4^2 \times 3^6\times10^2}{5^6}\]

OpenStudy (mertsj):

So pick up your calculator and multiply 4^2 by 3^6 and divide it by 5^6

OpenStudy (mertsj):

I got .746496. What did you get?

OpenStudy (anonymous):

.746?

OpenStudy (anonymous):

yeah me too

OpenStudy (mertsj):

Now it has to be multiplied by 10^2 which is 100.

OpenStudy (mertsj):

So the answer will be 74.6 or if it is rounded, it will be 75

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw